# High-dimensional covariance estimation by pairwise likelihood truncation

### v1

@@ Abstract
Pairwise likelihood offers a useful approximation to the full likelihood function for covariance estimation in high-dimensional context. It simplifies high-dimensional dependencies by combining marginal bivariate likelihood objects, thereby making estimation more manageable. In certain models, including the Gaussian model, both pairwise and full likelihoods are known to be maximized by the same parameter values, thus retaining optimal statistical efficiency, when the number of variables is fixed. Leveraging this insight, we introduce the estimation of sparse high-dimensional covariance matrices by maximizing a truncated version of the pairwise likelihood function, which focuses on pairwise terms corresponding to nonzero covariance elements. To achieve a meaningful truncation, we propose to minimize the discrepancy between pairwise and full likelihood scores plus a $L_{1}$-penalty discouraging the inclusion of uninformative terms. Differently from other regularization approaches, our method selects whole pairwise likelihood objects rather than individual covariance parameters, thus retaining the inherent unbiasedness of the pairwise likelihood estimating equations. This selection procedure is shown to have the selection consistency property as the covariance dimension increases exponentially fast. As a result, the implied pairwise likelihood estimator is consistent and converges to the oracle maximum likelihood estimator that assumes knowledge of nonzero covariance entries.
@@ 1 Introduction
Covariance estimation is a fundamental problem in multivariate analysis with countless applications in many fields of empirical research. For example, in genetics, it helps understand complex traits by identifying genetic correlations; in finance, it is instrumental for risk management and asset allocation; in environmental science and climate studies, it is used to study interactions between climate variables and predict future trends. In statistics, covariance estimation is fundamental for complex dependence models such as principal component analysis and factor models. With the emergence of new technology, covariance estimation has become increasingly challenging due to the growing number of variables involved. This growth makes traditional methods, such as maximum likelihood (ML) estimation, prone to error accumulation.
@@
maximizes $\ell({\bm{\theta}})$ and is the ML estimator of ${\bm{\Theta}}$. When the data are non-Gaussian or weakly dependent, $\ell({\bm{\theta}})$ is interpreted as a quasi-likelihood function and is still a valid objective function for estimation. The high-dimensional setting where the dimension $p$ is larger than the sample size $n$ is of particular interest in current applications but poses nontrivial challenges. In this setting, the ML estimator, without additional restrictions on the covariance structure, is singular and inconsistent due to the accumulation of excessive estimation error and too many free parameters (Bai and Yin, 1993; Bai et al., 1988). To deal with the overwhelming number of parameters, a common assumption in the literature also adopted here is that ${\bm{\Theta}}$ is sparse, meaning that a number of its entries are exactly zero. Particularly, we define the index set =${\mathcal{E}}=\{jk:j\leq k,\theta_{jk}\neq 0\}$ corresponding to diagonal elements of ${\bm{\Theta}}$ and the pairs for which covariances are actually different from zero. We use $m_{0}=\text{card}({\mathcal{E}})$ to denote the cardinality of ${\mathcal{E}}$, while =$\setminus{\mathcal{E}}=\{jk:j<k;\theta_{jk}=0\}$ denotes the complement of ${\mathcal{E}}$.
@@
Another popular approach to obtain sparse estimates is thresholding (Bickel and Levina, 2008a; Cai and Zhou, 2012), which modifies the sample covariance matrix by setting a portion of the estimated off-diagonal elements to zero. Thresholding reduces noise accumulation by avoiding the estimation of irrelevant off-diagonal elements of ${\bm{\Theta}}$. The Hard thresholding sets to zero the $jk$th covariance parameter if $|S_{jk}|\leq\psi$, for some threshold $\psi$, while adaptive thresholding (Cai and Liu, 2011) also takes into account heterogeneous estimation uncertainty. Particularly, the estimated off-diagonal elements are set to zero if $|S_{jk}|/\text{SE}_{jk}\leq\psi$, $j\neq k$, where $\text{SE}_{jk}$ is the standard error of the marginal estimate $S_{jk}$. Generalized thresholding extends these approaches by replacing hard thresholding operators with soft-thresholding operators (Rothman et al., 2009). Other established approaches include regularization on the eigenstructure of the matrix (Daniels and Kass, 2001; Ledoit and Wolf, 2012), banding and tapering (Bickel and Levina, 2008b; Bien, 2019; Bien et al., 2016; Furrer et al., 2006; Xue and Zou, 2014).
@@
Overall, PL estimators show considerable potential for high-dimensional covariance estimation due to simplicity in defining the objective function, their computational advantages and the statistical efficiency in common model families when $p$ is small compared to the number of observations. However, the application of traditional PL estimators in large problems is still hindered by their poor statistical efficiency in the $p>n$ setting. The main challenges are similar to those for ML estimation: without some form of restriction on the model structure, PL estimators still accumulate excessive error and are not generally consistent when the internal correlation among low-dimensional data subsets grows too fast as $p$ diverges (Cox and Reid, 2004; Ferrari et al., 2016; Huang and Ferrari, 2022). More recently, building on the success of shrinkage methods for the full likelihood, numerous studies have extended the use of sparsity-inducing penalties to the composite likelihood framework, applying it to various specialized models (Bradic et al., 2011; Xue et al., 2012; Gao and Carroll, 2017; Hui et al., 2018).
@@
Here we adopt a completely different approach compared to existing penalized procedures for PL estimation. Instead of shrinking individual elements of the parameter vector ${\bm{\theta}}$ to zero, we propose to discourage the inclusion of entire pairwise likelihood objects in the final PL objective function. To highlight this important feature, this approach will be referred to as truncated pairwise likelihood (TPL) estimation in the remainder of the paper. The main conceptual innovation lies in recognizing that these partial pairwise components are not strictly statistical parameters. Therefore, penalizing the log-pairwise likelihood function may not be the most suitable approach for eliminating such components, and instead, we opt for an alternative statistical criterion. Particularly, theory of unbiased estimating equations suggests that one should minimize the distance between the candidate PL and the ML score functions or, equivalently, maximize the statistical efficiency of the underlying PL estimator (Heyde, 1997; Lindsay et al., 2011). Based on this understanding, we propose to minimize a convex criterion related to statistical efficiency with the addition of a weighted $L_{1}$-penalty to discourage the selection of too many noisy or redundant pairwise scores terms.
@@
One advantage of the TPL method is that it inherently yields unbiased estimates. While conventional penalization techniques induce bias in the estimating equations to shrink estimates towards zero, our penalized criterion operates by selecting entire bivariate models, thereby preserving the unbiasedness from the corresponding pairwise score function, provided that the set of selected pairwise likelihoods is sufficiently comprehensive, i.e. it includes at least those covariances that are non-zero. Moreover, for Gaussian data, these estimates coincide with the Maximum Likelihood Estimates (MLE), thus exhibiting optimal efficiency. From a theoretical viewpoint, the procedure is shown to have the selection consistency property as $p$ grows exponentially in $n$ (Theorem 4). A meaningful consequence is that selection consistency implies that our estimator converges in probability to the oracle ML estimator, i.e. the ML estimator that assumes knowledge of the true covariance structure.
@@ 2.1 An efficiency criterion for pairwise likelihood truncation
In this section, we introduce an empirical criterion for selecting the relevant pairwise contributions. We adopt the PL estimation framework due to a crucial property: the PL function attains its maximum at the same parameter values as the full likelihood function for certain covariance models within the exponential family, such as the Gaussian distribution; see Mardia (Mardia et al., 2009). Hence, the resulting PL estimator achieves full efficiency, at least when the sparsity structure is known in advance. Leveraging the PL framework enables us to simplify the complexity of the full likelihood while retaining a substantial amount of statistical information. Next, we formulate our statistical criterion with the multivariate Gaussian distribution in mind, which is common in multivariate analysis, though the criterion remains applicable in other scenarios as well.
@@
with respect to ${\mathbf{w}}$, where ${\mathbf{u}}^{ML}({\bm{\theta}};{\bm{X}})=\partial\log f({\bm{X}};{\bm{\theta}})/\partial{\bm{\theta}}$ is the ML score and $f(\cdot;{\bm{\theta}})$ is the $p$-variate density of ${\bm{X}}$, which may be unknown. Minimizing (7) is interpreted as to decreasing the distance between ML and PL scores, effectively maximizing statistical efficiency for a given level of sparsity. Particularly, the weighted $L_{1}$-penalty $\lambda n^{-1}\sum_{j<k}|w_{jk}|\theta^{-2}_{jk}$ discourages the inclusion of too many pairwise contributions, due to its geometric properties; particularly, several elements of ${\mathbf{w}}$ become zero when $\lambda$ is sufficiently large. This penalty is adaptive in the sense that for $\theta_{jk}$ closer 00, the $jk$th pairwise score receives a larger penalty. Differently from the adaptive Lasso penalty of Zou Zou (2006) for sparse regression, here the penalty focuses on coefficients $w_{jk}$ associated with entire score terms, while the weight $\theta^{-2}_{jk}$ depends on the covariance parameters but not on ${\mathbf{w}}$, thus playing the role of a fixed weight for the $jk$th term.
Straightforward algebra shows that (7) can be re-written as a convex criterion, not depending on the maximum likelihood score ${\mathbf{u}}^{ML}$; hence a solution can be efficiently obtained using available algorithms. Particularly, ignoring an irrelevant term not depending on ${\mathbf{w}}$, we expand (7) and define the objective function:
@@ 2.2 Covariance tresholding by pairwise likelihood truncation
Denote $\hat{\bm{\theta}}$ the TPL estimator of ${\bm{\theta}}$, with elements $\hat{\theta}_{jk}$ estimating $\theta_{jk}$ in ${\bm{\theta}}$. In this section, we describe $\hat{\bm{\theta}}$ based on the criterion introduced in Section 2.1. Given an i.i.d. sample ${\bm{X}}^{(1)},\dots,{\bm{X}}^{(n)}$ from the distribution of ${\bm{X}}$, we solve the PL estimating equations $\sum_{i=1}^{n}{\mathbf{u}}(\hat{\bm{\theta}},{\mathbf{w}};{\bm{X}}^{(i)})=\bf 0$, with PL score ${\mathbf{u}}(\hat{\bm{\theta}},{\mathbf{w}};{\bm{X}}^{(i)})$ defined in (6). It is easy to check that the solution corresponds to $\hat{\theta}_{jj}=S_{jj}$, $1\leq j\leq p$, and $\hat{\theta}_{jk}=S_{jk}$ for those pairs $jk$ such that $w_{jk}\neq 0$ and this solution is invariant with respect to the elements of ${\mathbf{w}}$ different from zero. Particularly, with $u_{jj}$ defined in (2.1), the marginal equation
$\sum_{i=1}^{n}u_{jj}=\sum_{i=1}^{n}\frac{(X_{j}^{(i)})^{2}-\hat{\theta}_{jj}}{\hat{\theta}^{2}_{jj}}=0,$ (11)
is solved at $\hat{\theta}_{jj}=S_{jj}=n^{-1}\sum_{i=1}^{n}(X_{j}^{(i)})^{2}$, for all $1\leq j\leq p$. Substituting $\hat{\theta}_{jj}=S_{jj}$ and $\hat{\theta}_{kk}=S_{kk}$ into the pair-wise estimating equation $\sum_{i=1}^{n}{\mathbf{u}}_{jk}(\hat{\bm{\theta}};{\bm{X}}^{(i)})$ gives
$\sum_{i=1}^{n}u_{jk}^{[1]}=-\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\hat{\theta}_{jk}}{{S_{kk}}^{-1}\hat{\theta}_{jk}^{-1}(S_{jj}S_{kk}-\hat{\theta}_{jk}^{2})^{2}},\quad\sum_{i=1}^{n}u_{jk}^{[2]}=-\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\hat{\theta}_{jk}}{{S_{jj}}^{-1}\hat{\theta}_{jk}^{-1}(S_{jj}S_{kk}-\hat{\theta}_{jk}^{2})^{2}},$
$\sum_{i=1}^{n}u_{jk}^{[3]}=\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\hat{\theta}_{jk}}{(\hat{\theta}_{jk}^{2}+S_{jj}S_{kk})^{-1}(S_{jj}S_{kk}-\hat{\theta}_{jk}^{2})^{2}},$
where $u_{jk}^{[t]}$, $t=1,2,3$, are defined in (2.1). It is easy to see that the pairwise equations are solved at $\hat{\theta}_{jk}=S_{jk}=n^{-1}\sum_{i=1}^{n}X_{j}X_{k}$ as long as $w_{jk}\neq 0$. On the other hand, if $w_{jk}=0$, we set $\hat{\theta}_{jk}=0$, since the corresponding score contribution for estimating $\theta_{jk}$ is exacly zero.
A sparse covariance estimate $\hat{\bm{\Theta}}$ is then computed by applying the following thresholding mechanism to the entries of ${\mathbf{S}}$:
@@
Equation (13) represents a plug-in estimate of the population criterion $d_{\lambda}({\mathbf{w}};{\bm{\theta}})$ in Equation (8), obtained by setting ${\bm{\theta}}={\mathbf{s}}$ and with expectations replaced by averages. The threshold estimator in Equation (12) describes a simple mechanism whereby sparsity on ${\bm{\theta}}$ is induced by the relative information contained in the pairwise likelihood scores. Since ${\mathbf{w}}$ is not, strictly speaking, a statistical parameter, but it rather defines the structure of the estimating equations, selection of ${\mathbf{w}}$ should require special considerations. Particularly, our objective is completely different from that in classic regularization methods which obtain sparsity of ${\bm{\theta}}$ by maximizing the full or pairwise likelihood functions subject to a regularization constraint, such as the Lasso constraint $\sum_{j<k}|\theta_{jk}|<u$, $u>0$, on the parameter space.
@@
Positive definiteness of ${\bm{\Theta}}$ implies linear independence of elements in ${\bm{X}}$, ensuring that the score elements defined in (2.1) are nonzero with probability one. Thus by definition, $\hat{\mathbf{J}}_{\hat{\mathcal{E}}}$ is positive definite almost surely, leading to a unique solution of $\hat{\mathbf{w}}$, unlike in Lasso regression where columns of the design matrix are required to be in general position for similar results.
From Theorem 1 it follows that, for sufficiently large $\lambda>0$, a relatively small subset of scores is selected. Formula (21) suggests that the $jk$th score is selected when it contributes relatively large information in the overall pairwise likelihood relative to the size of $S_{jk}$. The extent of the contribution is measured by the difference between the marginal estimated Fisher information $\mathrm{diag}(\hat{\mathbf{J}})_{jk}$ for the $jk$th score and the information already present in other selected scores represented by the linear combination $\hat{\mathbf{J}}_{jk,\cdot}\hat{\mathbf{w}}$.

### v2

@@ Abstract
Pairwise likelihood is a useful approximation to the full likelihood function for covariance estimation in high-dimensional context. It simplifies high-dimensional dependencies by combining marginal bivariate likelihood objects, thus making estimation more manageable. In certain models, including the Gaussian model, both pairwise and full likelihoods are maximized by the same parameter values, thus retaining optimal statistical efficiency, when the number of variables is fixed. Leveraging on this insight, we introduce estimation of sparse high-dimensional covariance matrices by maximizing a truncated version of the pairwise likelihood function, obtained by including pairwise terms corresponding to nonzero covariance elements. To achieve a meaningful truncation, we propose to minimize the $L_{2}$-distance between pairwise and full likelihood scores plus an $L_{1}$-penalty discouraging the inclusion of uninformative terms. Differently from other regularization approaches, our method focuses on selecting whole pairwise likelihood objects rather than shrinking individual covariance parameters, thus retaining the inherent unbiasedness of the pairwise likelihood estimating equations. This selection procedure is shown to have the selection consistency property as the covariance dimension increases exponentially fast. Consequently, the implied pairwise likelihood estimator is consistent and converges to the oracle maximum likelihood estimator assuming knowledge of nonzero covariance entries.
@@ 1 Introduction
Covariance estimation is a fundamental problem in multivariate analysis and has a range of applications in many scientific fields. In genetics and genomics, it helps our understanding of complex traits by identifying genetic correlations. In finance, it is essential for risk management and asset allocation strategies. In environmental science and climate studies, it helps us represent interactions between climate variables and predicting future trends. In statistics, covariance estimation is a cornerstone for more complex dependence models such as principal component analysis, factor analysis, and linear models. As new technology emerges, covariance estimation becomes increasingly challenging due to the growth of the number of variables involved, leading traditional methods to become more computationally intensive and susceptible to statistical error accumulation.
@@
maximizes $\ell({\bm{\theta}})$ and is the ML estimator of ${\bm{\Theta}}$. When the data are non-Gaussian or weakly dependent, $\ell({\bm{\theta}})$ is interpreted as a quasi-likelihood function, yet it remains a valid objective function for estimation. The high-dimensional setting where the dimension $p$ is larger than the sample size $n$ is of particular interest in current applications, but poses nontrival challenges. In this setting, the ML estimator, without additional restrictions on the covariance structure, is singular and inconsistent due to the accumulation of excessive estimation error and too many free parameters (Bai and Yin, 1993; Bai et al., 1988). To deal with the overwhelming number of parameters, a common assumption in the literature also adopted here is that ${\bm{\Theta}}$ is sparse, meaning that a number of its entries are exactly zero. Particularly, we define the index set =${\mathcal{E}}=\{jk:j\leq k,\theta_{jk}\neq 0\}$ corresponding to diagonal elements of ${\bm{\Theta}}$ and the pairs for which covariances are actually different from zero. We use $m_{0}=\text{card}({\mathcal{E}})$ to denote the cardinality of ${\mathcal{E}}$, while =$\setminus{\mathcal{E}}=\{jk:j<k;\theta_{jk}=0\}$ denotes the complement of ${\mathcal{E}}$.
@@
Another popular approach to obtain sparse estimates is thresholding (Bickel and Levina, 2008a; Cai and Zhou, 2012), which modifies the sample covariance matrix by setting a portion of the estimated off-diagonal elements to zero. Thresholding reduces noise accumulation by avoiding the estimation of irrelevant off-diagonal elements of ${\bm{\Theta}}$. Hard thresholding sets to zero the $jk$th covariance parameter if $|S_{jk}|\leq\psi$, for some threshold $\psi>0$. Adaptive thresholding (Cai and Liu, 2011) also takes into account heterogeneous estimation uncertainty; particularly, the estimated off-diagonal elements are set to zero if $|S_{jk}|/\text{SE}_{jk}\leq\psi$, $j\neq k$, where $\text{SE}_{jk}$ is the standard error of the marginal estimate $S_{jk}$. Generalized thresholding extends these approaches by replacing hard thresholding operators with soft-thresholding operators (Rothman et al., 2009). Other established methods for high-dimensional covariance estimation include regularization on the eigenstructure of the matrix (Daniels and Kass, 2001; Ledoit and Wolf, 2012), banding and tapering (Bickel and Levina, 2008b; Bien, 2019; Bien et al., 2016; Furrer et al., 2006; Xue and Zou, 2014).
@@
Overall, PL estimators show considerable potential for high-dimensional covariance estimation due to the simplicity in defining the objective function, their computational advantages and the statistical efficiency in common model families when $p$ is small compared to the number of observations. However, the application of traditional PL estimators in large problems is still hindered by their poor statistical efficiency in the $p>n$ setting. The main challenges are similar to those for ML estimation: without some form of restriction on the model structure, PL estimators accumulate excessive error and are not generally consistent as $p$ diverges (Cox and Reid, 2004; Ferrari et al., 2016; Huang and Ferrari, 2022). More recently, building on the success of shrinkage methods for the full likelihood, numerous studies have extended the use of sparsity-inducing penalties to the composite likelihood framework, applying it to various specialized models (Bradic et al., 2011; Xue et al., 2012; Gao and Carroll, 2017; Hui et al., 2018).
@@
Here we adopt a completely different approach compared to existing penalized procedures for PL estimation. Instead of shrinking individual elements of the parameter vector ${\bm{\theta}}$ to zero, we propose to discourage the inclusion of entire pairwise likelihood objects in the final PL objective function. To highlight this important feature, this approach will be referred to as truncated pairwise likelihood (TPL) estimation in the remainder of the paper. The main conceptual innovation lies in recognizing that these partial pairwise components are not strictly statistical parameters. Therefore, penalizing the log-pairwise likelihood function may not be the most suitable approach for eliminating such components, and instead we opt for an alternative statistical criterion. Particularly, the theory of unbiased estimating equations suggests that one should minimize the distance between the candidate PL and the ML score functions or, equivalently, maximize the statistical efficiency of the underlying PL estimator (Heyde, 1997; Lindsay et al., 2011). Based on this understanding, we propose to minimize a convex criterion related to statistical efficiency with the addition of a weighted $L_{1}$-penalty to discourage the selection of too many noisy or redundant pairwise scores terms.
@@
One advantage of the TPL method is that it inherently yields unbiased estimates. While conventional penalization techniques introduce bias in the estimating equations to shrink estimates towards zero, our penalized criterion operates by selecting entire bivariate models, thereby preserving the unbiasedness from the corresponding pairwise score function, provided that the set of selected pairwise likelihoods is sufficiently comprehensive, i.e., it includes at least those covariances that are non-zero. Moreover, for Gaussian data, these estimates coincide with the Maximum Likelihood Estimates (MLE), thus yielding optimal efficiency. From a theoretical viewpoint, the procedure is shown to have the selection consistency property as $p$ grows exponentially in $n$ (Theorem 4). A meaningful consequence is that selection consistency implies that our estimator converges in probability to the oracle ML estimator, i.e., the ML estimator that assumes knowledge of the true covariance structure.
@@ 2.1 An efficiency criterion for pairwise likelihood truncation
In this section, we introduce an empirical criterion for selecting the relevant pairwise contributions. We adopt the PL estimation framework due to a crucial property: the PL function attains its maximum at the same parameter values as the full likelihood function for certain covariance models within the exponential family, such as the Gaussian distribution; see Mardia (Mardia et al., 2009). Hence, the resulting PL estimator achieves full efficiency, at least when the sparsity structure is known in advance. Leveraging the PL framework enables us to simplify the complexity of the full likelihood while retaining a substantial amount of statistical information. Next, we formulate our statistical criterion with the multivariate Gaussian distribution in mind, which is common in multivariate analysis, but the criterion remains applicable in other scenarios as well.
@@
with respect to ${\mathbf{w}}$, where ${\mathbf{u}}^{ML}({\bm{\theta}};{\bm{X}})=\partial\log f({\bm{X}};{\bm{\theta}})/\partial{\bm{\theta}}$ is the ML score and $f(\cdot;{\bm{\theta}})$ is the $p$-variate density of ${\bm{X}}$, which may be unknown. Minimizing (7) is interpreted as decreasing the distance between ML and PL scores, effectively maximizing statistical efficiency for a given level of sparsity. Particularly, the weighted $L_{1}$-penalty $\lambda n^{-1}\sum_{j<k}|w_{jk}|\theta^{-2}_{jk}$ discourages the inclusion of too many pairwise contributions, due to its geometric properties and several elements of ${\mathbf{w}}$ become zero when $\lambda$ is sufficiently large. This penalty is adaptive in the sense that for $\theta_{jk}$ closer to zero, the $jk$th pairwise score receives a larger penalty. Differently from the adaptive Lasso penalty of Zou Zou (2006) for sparse regression, here the penalty focuses on coefficients $w_{jk}$ associated with entire score terms, while the quantity $\theta^{-2}_{jk}$ depends on the covariance parameters but not on ${\mathbf{w}}$, thus playing the role of a fixed weight for the $jk$th term.
Straightforward algebra shows that (7) can be re-written as convex criterion not depending on the maximum likelihood score ${\mathbf{u}}^{ML}$; hence a solution can be efficiently obtained using available algorithms. Particularly, ignoring an irrelevant term not depending on ${\mathbf{w}}$, we expand (7) and define the objective function:
@@ 2.2 Covariance tresholding by pairwise likelihood truncation
In this section, we describe our approach to obtain a sparse estimate of ${\bm{\theta}}$. Given an i.i.d. sample ${\bm{X}}^{(1)},\dots,{\bm{X}}^{(n)}$ from the distribution of ${\bm{X}}$, we solve the estimating equations $\sum_{i=1}^{n}{\mathbf{u}}({\bm{\theta}},{\mathbf{w}};{\bm{X}}^{(i)})=\bf 0$. It is easy to check that the solution corresponds to $\theta_{jj}=S_{jj}$, $1\leq j\leq p$, and $\theta_{jk}=S_{jk}$ for those pairs $jk$ such that $w_{jk}\neq 0$. This solution is invariant with respect to the elements of ${\mathbf{w}}$ different from zero. Particularly, the marginal estimating equation
$\sum_{i=1}^{n}u_{jj}({\bm{\theta}};{\bm{X}}^{(i)})=\sum_{i=1}^{n}w_{jj}\left(\dfrac{(X_{j}^{(i)})^{2}-\theta_{jj}}{\theta^{2}_{jj}}\right)=0,$ (11)
is solved by $\theta_{jj}=S_{jj}=n^{-1}\sum_{i=1}^{n}(X_{j}^{(i)})^{2}$, for all $1\leq j\leq p$. Substituting $\theta_{jj}=S_{jj}$ and $\theta_{kk}=S_{kk}$ into $\sum_{i=1}^{n}{\mathbf{u}}_{jk}({\bm{\theta}};{\bm{X}}^{(i)})$ gives
$\sum_{i=1}^{n}u_{jk}^{[1]}=-\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\theta_{jk}}{{S_{kk}}^{-1}\theta_{jk}^{-1}(S_{jj}S_{kk}-\theta_{jk}^{2})^{2}},\quad\sum_{i=1}^{n}u_{jk}^{[2]}=-\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\theta_{jk}}{{S_{jj}}^{-1}\theta_{jk}^{-1}(S_{jj}S_{kk}-\theta_{jk}^{2})^{2}},$
$\sum_{i=1}^{n}u_{jk}^{[3]}=-\frac{\sum_{i=1}^{n}X_{j}^{(i)}X_{k}^{(i)}-n\theta_{jk}}{(\theta_{jk}^{2}+S_{jj}S_{kk})^{-1}(S_{jj}S_{kk}-\theta_{jk}^{2})^{2}},$
where $u_{jk}^{[r]}$, $r=1,2,3$, are defined in (2.1). It is easy to see that the estimating equations are solved by $\theta_{jk}=S_{jk}=n^{-1}\sum_{i=1}^{n}X_{j}X_{k}/n$ as long as $w_{jk}\neq 0$. On the other hand, if $w_{jk}=0$, we set $\theta_{jk}=0$, since the corresponding score contribution for estimating $\theta_{jk}$ is exacly zero.
The TPL estimator $\hat{\bm{\theta}}$ of ${\bm{\theta}}$ is then obtained by applying the following thresholding mechanism to the entries of ${\mathbf{S}}$:
@@
Equation (13) represents a plug-in estimate of the population criterion $d_{\lambda}({\mathbf{w}};{\bm{\theta}})$ in Equation (8), obtained by setting ${\bm{\theta}}={\mathbf{s}}$ and with expectations replaced by averages. The threshold estimator in Equation (12) describes a simple mechanism whereby sparsity on ${\bm{\theta}}$ is induced by the relative information contained in the pairwise likelihood scores. Since ${\mathbf{w}}$ is not, strictly speaking, a statistical parameter, but rather defines the structure of the estimating equations, selection of ${\mathbf{w}}$ should require special considerations. Notably, our objective is different from that in classic regularization methods which obtain sparsity of ${\bm{\theta}}$ by maximizing the full or pairwise likelihood functions subject to a regularization constraint, such as the Lasso constraint $\sum_{j<k}|\theta_{jk}|<u$, $u>0$, on the parameter space.
@@
Positive definiteness of ${\bm{\Theta}}$ implies linear independence of elements in ${\bm{X}}$, ensuring that the score elements defined in (2.1) are nonzero with probability one. Thus by definition, $\hat{\mathbf{J}}_{\hat{\mathcal{E}}}$ is positive definite almost surely, leading to a unique solution of $\hat{\mathbf{w}}$; in contrast, in Lasso regression the columns of the design matrix are required to be in general position to yield a similar result.
From Theorem 1 it follows that, for sufficiently large $\lambda>0$, a relatively small subset of scores is selected. Formula (21) suggests that the $jk$th score is selected when it contributes relatively large information in the overall pairwise likelihood relative to the size of $S_{jk}$. The extent of the contribution is measured by the difference between the marginal estimated Fisher information for the $jk$th score, $\mathrm{diag}(\hat{\mathbf{J}})_{jk}$, and the information already present in other selected scores, represented by the linear combination $\hat{\mathbf{J}}_{jk,\cdot}\hat{\mathbf{w}}$.